OpenStudy (anonymous):

Using a directrix of y = -3 and a focus of (2, 1), what quadratic function is created?

OpenStudy (anonymous):

you need the vertex first do you know it ?

OpenStudy (anonymous):

how can you find that

OpenStudy (anonymous):



OpenStudy (anonymous):

it is half way between the directrix and the focus

OpenStudy (anonymous):

half way between \((2,1)\) and \((2,-3)\) is \((2,-1)\)

OpenStudy (anonymous):

ohh okay gotcha

OpenStudy (anonymous):

and the picture tells you it opens up. so the \(x\) term is squared \[4p(y-k)=(x-h)^2\]

OpenStudy (anonymous):

you have \((h,k)\) is \((2,-1)\) so far we have \[4p(y+1)=(x-2)^2\] all we need is \(p\)

OpenStudy (anonymous):

\(p\) is the distance between the focus and the vertex the distance between \((2,-1)\) and \((2,1)\) is \(2\) so \(p=2\)

OpenStudy (anonymous):

giving \[8(y+1)=(x-2)^2\]

OpenStudy (anonymous):

and as usual we can cheat and check the answer http://www.wolframalpha.com/input/?i=parabola+8%28y%2B1%29%3D%28x-2%29^2

OpenStudy (anonymous):

ooh you need to solve this for \(y\) don't you?

OpenStudy (anonymous):

i think so

OpenStudy (anonymous):

we cheat for that too http://www.wolframalpha.com/input/?i=+8%28y%2B1%29%3D%28x-2%29^2

OpenStudy (anonymous):

not that it is that difficult to solve this for \(y\) looks like you get y = x^2/8-x/2-1/2

OpenStudy (anonymous):

hope the steps are clear ish

OpenStudy (anonymous):

yes i got it thanks again

OpenStudy (anonymous):